### v1

# VideoGen-of-Thought: A Collaborative Framework for Multi-Shot Video Generation

Abstract
Current video generation models excel at generating short clips but still struggle with creating multi-shot, movie-like videos. Existing models trained on large-scale data on the back of rich computational resources are unsurprisingly inadequate for maintaining a logical storyline and visual consistency across multiple shots of a cohesive script since they are often trained with a single-shot objective. To this end, we propose VideoGen-of-Thought (VGoT), a collaborative and training-free architecture designed specifically for multi-shot video generation. VGoT is designed with three goals in mind as follows. Multi-Shot Video Generation: We divide the video generation process into a structured, modular sequence, including (1) Script Generation, which translates a curt story into detailed prompts for each shot; (2) Keyframe Generation, responsible for creating visually consistent keyframes faithful to character portrayals; and (3) Shot-Level Video Generation, which transforms information from scripts and keyframes into shots; (4) Smoothing Mechanism that ensures a consistent multi-shot output. Reasonable Narrative Design: Inspired by cinematic scriptwriting, our prompt generation approach spans five key domains, ensuring logical consistency, character development, and narrative flow across the entire video. Cross-Shot Consistency: We ensure temporal and identity consistency by leveraging identity-preserving (IP) embeddings across shots, which are automatically created from the narrative. Additionally, we incorporate a cross-shot smoothing mechanism, which integrates a reset boundary that effectively combines latent features from adjacent shots, resulting in smooth transitions and maintaining visual coherence throughout the video. Our experiments demonstrate that VGoT surpasses existing video generation methods in producing high-quality, coherent, multi-shot videos. Project page: https://cheliosoops.github.io/VGoT
Figure 1: Illustration of VideoGen-of-Thought (VGoT). (a) Comparison of existing methods with VGoT in multi-shot video generation. Existing methods struggle with maintaining consistency and logical coherence across multiple shots, while VGoT effectively addresses these challenges through a multi-shot generation approach. (b) Overview of our proposed framework VGoT, which is consist of the Script Module which generates detailed shot descriptions from five domains, the KeyFrame Module to create keyframes from scripts, the Shot-Level Video Module which synthesizes video latents on conditional with keyframes and scripts, and the Smooth Module ensures seamless transitions across shots, resulting in a cohesive video narrative.
1 Introduction
Recent advancements in video generation techniques have yielded impressive results, particularly in creating short, visually appealing clips [6, 8, 9, 69, 28]. These advancements have been powered by increasingly sophisticated generative models, ranging from diffusion models [29, 56, 49, 6] to auto-regressive models [20, 65, 39, 62], supported by large-scale datasets [31, 50, 51]. These methods have enabled the generation of high-quality, realistic short videos. However, generating multi-shot videos from a brief user input script remains a substantial challenge. Despite existing works successfully achieving long-term video generation, the primary focus has been on extending the duration of single-shot videos [19, 40, 61, 35]. In contrast, multi-shot video generation involves additional complexities in ensuring coherence across different shots. Moreover, while simply scaling up models or using larger datasets, as done in single-shot long video generation methods [63, 45], can enhance the visual quality of individual clips, it requires vast computational resources and extensive training times. More importantly, they fail to solve the challenges inherent in multi-shot video generation, which is a more complex task that requires maintaining reasonable logic in the narrative, and visual consistency across different scenes—all of which are essential for producing multi-shot, movie-like content.
The limitations of existing approaches underscore a fundamental gap in the field: increasing model size and dataset volume alone is insufficient for generating reasonable and consistent videos with multiple shots. The primary challenge is not solely to achieve high visual fidelity but also to ensure that each shot seamlessly connects to the next, thereby contributing to a coherent and compelling narrative structure. When applied to multi-shot video generation, we observed that current methods often fall short in maintaining a consistent storyline and visual dynamics across multiple shots [8, 9, 69]. The absence of a structured approach for organizing and managing the generation process often results in a breakdown of narrative flow, with logical transitions between scenes being lost. Therefore, there is a pressing need for an approach that effectively balances visual quality with narrative consistency, managing the generation process without relying purely on computational scaling.
Inspired by the “thought process” behind the creation of real movies, we propose VideoGen-of-Thought (VGoT)—a collaborative framework designed to tackle these challenges and create multi-shot videos in a step-by-step manner. An existing approach with a similar manner called “Chain of Thought” [64] has been proven efficient in Large Language Models (LLMs) [1, 15] and some visual perception tasks [18]. In this paper, we extend this approach to video generation by breaking down the complex process into four distinct yet collaborative modules as shown in Fig. 1: Script Generation Module converts an simple and high-level story description into detailed prompts for each shot to provide a solid narrative foundation; Keyframe Generation Module generates visually consistent keyframes based on the script, ensuring consistency in the appearance of characters, environments, and objects; Shot-Level Video Module produces video latents from corresponding keyframes and prompts at each shot that represents a shot-level video clip; finally, the Smooth Module integrates a cross-shot smoothing mechanism that processes information embeddings for the upcoming shot when the current one is nearing completion, ensuring a seamless connection facilitated by a reset boundary. This mechanism provides a smooth transition between shots, resulting in an end-to-end shot-by-shot video generation process that maintains visual and narrative continuity across the entire video.
As a text-to-video [53, 57] (T2V) architecture, our method VGoT employs a structured prompt design approach inspired by cinematic scriptwriting to ensure the generated videos are reasonable. Given a one-sentence script as user input, our method first prepares it into short descriptions for each shot and then extends them into five domains including character, background, relations, camera pose, and High Dynamic Range (HDR) descriptions—which together form a comprehensive movie-like script, as shown in the left part of Fig 2. Reasonability here refers to the logical flow and narrative coherence of the generated video, ensuring that each shot contributes meaningfully to the storyline. Furthermore, our method also tackles the issue of consistency across shots by introducing mechanisms for both temporal and visual coherence. Specifically, we use identity-preserving (IP) embeddings from avatars automatically derived from the narrative to maintain consistent character portrayal throughout the video. Additionally, our cross-shot smoothing mechanism, which leverages latent features from adjacent shots, ensures seamless transitions, resulting in a cohesive and consistent multi-shot video.
Our experimental results demonstrate the efficacy of VGoT in generating high-quality, logically consistent, multi-shot videos. We evaluated VGoT using ten different stories, and a fair comparison with existing baselines highlights the superior performance of our approach in terms of narrative coherence, visual quality, and cross-shot consistency. The main contributions of this paper are as follows:
  1. 1.
    We propose a collaborative, training-free architecture VideoGen-of-Thought for multi-shot video generation, addressing the key challenges of reasonability and consistency across multiple shots.
  2. 2.
    We introduce a structured prompt design that emulates cinematic scriptwriting, ensuring logical coherence and narrative depth in generated videos.
  3. 3.
    We develop a cross-shot consistency mechanism using identity-preserving embeddings and a cross-shot smoothing technique to maintain visual continuity and temporal coherence across different shots.
2 Related Work
Video generation has progressed significantly following the great success of diffusion models, leading to two important research categories: long video synthesis and multi-shot story generation. These areas focus on generating high-quality, consistent videos either as extended single shots or as coherent sequences across multiple scenes.
Long Video Synthesis. Long video synthesis has advanced through diffusion-based methods and autoregressive approaches. Diffusion models based on the Latent Diffusion Model [49] utilize iterative refinement to generate visually consistent frames and have been effective for short sequences and other domains [26, 6, 8, 9, 68, 66, 7, 25, 71, 79, 22, 32, 44, 16, 76, 54, 78, 82, 46, 41, 42, 58, 80, 43, 38]. However, they struggle with maintaining coherence over extended video lengths. Autoregressive methods [20, 36, 73] predict frames sequentially but often face error accumulation, making long-term consistency challenging and computationally expensive. Additionally, training-free methods like FIFO-Diffusion [33] generate long sequences without training but lack mechanisms to manage transitions across shots, limiting their effectiveness in narrative-driven content. Overall, while these approaches achieve visual fidelity, they fail to ensure logical coherence across extended sequences. In contrast, VideoGen-of-Thought (VGOT) leverages a modular approach that includes cross-shot smoothing mechanisms to ensure both visual consistency and narrative coherence, offering a more holistic solution for generating long-form videos.
Multiple Shot Story Generation. Multiple Shot Story Generation focuses on maintaining narrative coherency across distinct scenes. Animate-a-Story [27] uses retrieval-augmented generation to ensure visual consistency but struggles with maintaining logical narrative transitions. DreamFactory [67] employs a multi-agent system for collaboration across shots, but lacks robust mechanisms for logical continuity, leading to inconsistencies. MovieDreamer [81] uses a hierarchical framework to manage coherence but relies heavily on predefined structures, limiting adaptability to dynamic stories. Flexifilm [40] attempts to address scene adaptability by incorporating flexible conditions, but it still falls short in ensuring logical cohesion between shots and often requires manual adjustments for consistency. Collectively, these methods often require significant manual intervention to achieve a consistent narrative, limiting scalability for fully automated storytelling. In contrast, our work provides a detailed and integrated framework that not only ensures cross-shot coherence through a novel smoothing mechanism but also employs rigorous evaluation methods to achieve both logical consistency and scalability, setting a new paradigm for multi-shot video generation.
3 Preliminaries and Problem Formulation
Figure 2: The FlowChart of VideoGen-of-Thought. Left: Shot descriptions are generated based on user prompts, describing various attributes including character details, background, relations, camera pose, and lighting HDR. Pre-shot descriptions provide a broader context for the upcoming scenes. Middle Top: Keyframes are generated using a text-to-image diffusion model conditioned with identity-preserving (IP) embeddings, which ensures consistent representation of characters throughout the shots. IP portraits help maintain visual identity consistency. Right: The shot-level video clips are generated from keyframes, followed by shot-by-shot smooth inference to ensure temporal consistency across different shots. This collaborative framework ultimately produces a cohesive narrative-driven video.
3.1 Preliminaries
Diffusion Models [29, 55, 56] are generative models that trained to approximate data distribution $p(x)$ from random noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. In the forward process, noise scheduled by the variance $\beta_{t}$ is gradually added as $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$ over $T$ time steps, resulting in noisy latent variables $x_{T}$. Then a learnable model $\mu_{\theta}$ parameterized by $\theta$ aims to denoise $x_{T}$ back to the original data sample $x_{0}$, with a parameterized distribution given by $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ where $\mu_{\theta}$ and $\Sigma_{\theta}$ are predicted mean and variance. The objective is to minimize the difference between predicted and true noise through the loss function:
$\mathcal{L}_{\text{uncond diff}}=\mathbb{E}_{x_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I}),t}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right]$ (1)
In the reverse process, an overall sample queue is maintained as $[x_{T};x_{T-1};\dots;x_{0}]$ at the pixel level. Latent diffusion models [49] successfully map this process into a lower-dimensional latent space using a variational autoencoder (VAE) [34, 70, 10, 37]. This transformation significantly enhances flexibility, allowing additional conditions to be incorporated through cross-attention mechanisms [60], which can be formulated as:
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (2)
where $Q$ is derived from the noisy image embeddings $e_{i}$, and $K$ and $V$ are derived from the text embeddings $e_{t}$ in text-to-image tasks. This mechanism has been further extended to the temporal dimension in conditional diffusion models for video generation [26, 6, 8, 74, 28, 75, 77]. For a video $V\in\mathbb{R}^{F\times H\times W\times 3}$ with $F$ frames and spatial resolution $H\times W$, a series of latent codes $\{z^{f}\}_{f=0}^{F-1}$, where $z^{f}\in\mathbb{R}^{h\times w\times d}$ with $d$ channels will be denoise in the same level. These video diffusion models, denoted as $\epsilon_{\theta}(z_{t},t,c)$, are capable of predicting $\hat{z}$, which can be decoded into video frames based on given conditions $c$.
While most existing methods aim to achieve long video generation through enlarging the shape of latent code $z$ and the size of model parameters $\theta$, FIFO-Diffusion [33] provides a new approach through denoising a queue of latents $Q_{k}=\{z^{f}\}_{f_{k}}^{f_{k}+n}$, updating $Q$ with DDIM[55] sampler $\Phi$ as:
$Q_{k}\leftarrow\Phi(Q_{k},\tau_{k},c;\epsilon_{\theta})$ (3)
for each timestep $\tau_{k}\in[0,T-n]$ from different level, in effect implementing long-term frame-by-frame prediction. Despite FIFO excelling in frame-by-frame prediction, it fails to handle the quick transition for different shots in multi-shot video generation. Moreover, as a text guided inference method, FIFO can’t receive conditions like image. In our work, we use the generative capability of diffusion models and process cross-shot smoothness with a FIFO-like mechanism at the shot level.
3.2 Problem Definition
Given an one-sentence user input $S$ containing $N$ number of shots, e.g., A set of one-sentence prompts, 30 shots, describe a story of a classic American woman Mary’s life, from birth to death, our method aims to generate a multi-shot, reasonable, and consistent video $V$ with minimal manual intervention. The key challenges are defined as:
  • Multi-Shot: Generating minute-level video containing plenty of shots.
  • Reasonability: Ensuring the generated video maintains a logical narrative and storyline.
  • Consistency: Maintaining temporal and identity consistency across different shots.
Our proposed solution, VideoGen-of-Thought (VGoT), starts from preparing the script $\left\{p_{i}\right\}^{N-1}_{i=0}$ for each shot, drafting keyframes $\left\{I_{i}\right\}^{N-1}_{i=0}$ with identity-preserving, then generating video latents $\{z^{f}\}_{f_{i}=0}^{F-1}$ with a cross-shot smooth mechanism within $N$ shots. VGoT provides a novel approach to overcoming the limitations of existing multiple shots video generation methods.
4 Method: VideoGen-of-Thought
In this section, we introduce each module in our proposed framework, VideoGen-of-Thought (VGoT), including Script Generation Module, Keyframe Generation Module, Shot-Level Video Generation Module, and Cross-Shot Smooth Module. By breaking down the complex task of long video generation into a series of smaller, well-defined modules, VGoT systematically addresses the key challenges of video length, logic, and cross-shot consistency. This modular decomposition allows each component to focus on specific aspects of the problem, resulting in long, reasonable, and consistent video results.
4.1 Module 1: Script Generation
The script generation task is modeled as a sequence generation problem using a pretrained large language model (LLM) [1, 15, 4]. The process starts with a short one-sentence user prompt $S$ with the number of shots $N$ defined, e.g., a set of one-sentence prompts, 30 shots, describe a story about three avatars (two friends Mike and Jane, one bad guy Tom) and their Fantasy Adventure in an ancient pyramid. We leverage an LLM to prepare short shot descriptions $S^{\prime}=\left\{s_{i}\right\}_{i=0}^{N-1}$ for each shot, which builds the foundation logic for the whole story. The script $p_{i}$ for one shot can be fulfilled from $s_{i}$ as:
$p_{i}(p_{cha},p_{b},p_{r},p_{cam},p_{h})=\mathcal{M}_{\text{LLM}}(s_{i},i)$ (4)
where $p_{cha}$ represents the character description, $p_{b}$ is the background description, $p_{r}$ describes the relation between characters and elements in the scene, $p_{cam}$ describes the camera pose, and $p_{h}$ defines the HDR lighting details. An example is shown in the left top of Figure 3. The LLM is conditioned on both the current $s_{i}$ and the previously generated prompts $p_{i-1}$ to maintain narrative coherence and avoid abrupt transitions between scenes. The algorithm for script generation is iterative, producing a prompt for each video shot as shown in Algorithm 1.
Algorithm 1 Script Generation for Long Video
1:User input $S$, total shot number $N$
2:Initialize $P\leftarrow\emptyset$, $p_{tmp}\leftarrow\emptyset$
3:Initialize Story $S^{\prime}=\{s_{1},s_{2},\dots,s_{N}\}$ with $S$
4:for each short shot description $s_{i}\in S$ do
5:     $p_{i}\leftarrow\mathcal{M}_{\text{LLM}}(s_{i},i)$
6:     Append $p_{i}$ to $P$
7:     $p_{tmp}\leftarrow p_{i}$
8:end for
9:return Prompt Sequence $P=\{p_{1},p_{2},\dots,p_{N}\}$
4.2 Module 2: Keyframe Generation
The keyframes generation task is based on the scripts prepared in the Script Module. For each shot, we use a pre-trained CLIP [47] encoder $E_{\text{text}}$ to obtain text embedding $e^{T}_{i}=E_{\text{text}}(p_{i})$. We leverage the knowledge of LLM used in script module to create some avatar descriptions for this story based on short shot descriptions $S^{\prime}$. For instance, given the $S^{\prime}$ of Mary’s life provided in Sec 3.2, LLM provides the avatar prompts $P_{a}=[P_{a1},P_{a2},\dots]$ of Child Mary, Teenager Mary, Mid-aged Mary, Elder-aged Mary, and Old Mary from five domains defined in Eq 4, and specify each shot prompt with a corresponding avatar.
As introduced in section 3, we use pretrained text-to-image diffusion models $\mathcal{M}_{I}$ [49, 59] to generate $I_{a}=[\mathcal{M}_{I}(P_{a1}),\mathcal{M}_{I}(P_{a2}),\dots]$ representing the avatar appearance from $P_{a}$ through cross-attention [60]. Then we employ IP-Adapter [72] to decouple the key $K$ and value $V$ in the cross-attention in $\mathcal{M}_{I}$ by introducing an additional block that involves the identity-preserving embedding from $I_{a}$. This mechanism ensures that each generated keyframe not only matches the given prompt but also maintains visual consistency with the given avatar through copied $K^{\prime}$ and $V^{\prime}$. The generation of each keyframe can be formalized as:
$I_{i}=\mathcal{M}_{I}(e^{T}_{i},e^{I}_{j})$ (5)
where $e^{T}_{i}$ is the text embedding from $p_{i}$ as mentioned above, $e^{I}_{j}$ is the image embedding of the $j$th IP avatar selected for $i$th shot which is obtained by a CLIP vision model [47]. This approach ensures keyframes $\left\{I_{i}\right\}^{N-1}_{i=0}$ maintain visual consistency in the visual appearance of characters and backgrounds style for the same avatar in different shots.
Figure 3: Visual comparison of VGoT and baselines
Figure 4: Visual showcases of VGoT generated multi-shot videos.
4.3 Module 3: Shot-level Video Generation
Through Script Module and KeyFrame Module, we have already got the scripts and keyframes required by each shot-level video. We employ DynamiCrafter [68] as a conditional video model $\mathcal{M}_{V}$ that is capable of taking both the keyframes and the corresponding textual prompts as input to generate a series of latents $Z_{i}=[z_{1},z_{2},...,z_{k}]$ representing the $i$th shot video clip $V_{i}$ which includes $k$ frames, and this process can be formulated as:
$Z_{i}=\mathcal{M}_{\text{V}}(e^{T}_{i},e^{I}_{i},\epsilon),$ (6)
where the $e^{T}_{i}$ and $e^{I}_{i}$ are the text embedding and image embedding of shot $i$, $Z_{i}\in\mathbb{R}^{f\times h\times w\times d}$ is a series of latent code that can be decoded into the video sequence $V_{i}$ containing $f$ frames generated for shot $i$, and $\epsilon_{i}$ is a noise map with same shape to initialize $Z_{i}$ Note that we use short shot description $s_{i}$ but not the detailed script $p_{i}$ to obtain the $e^{T}_{i}$ , because we found in our experiments that shot-level videos $V_{i}$ will be less dynamic when given prompts that are too complex even though we obtain much better keyframe through $p_{i}$. This phenomenon may have been caused by our video model $\mathcal{M}_{\text{V}}$ receiving simpler textual descriptions during training. As future work, this may be improved if $\mathcal{M}_{\text{V}}$ is trained with detailed scripts like Eq 4.
By incorporating both the keyframe and the textual description to obtain latents $z$ in this video generation process, we secure the generation of high-quality, smooth video clips in the subsequent module.
Table 1: Quantitative comparison with state-of-the-art T2V baselines. We compare average CLIP scores from five domains defined in Eq 4, and the average FC and SC scores within and across shots between VGoT and baseline models. We use $\mathbb{bold}$ to highlight the highest and $\underline{underline}$ for the second high.
Model CLIP (pcha) $\uparrow$ CLIP (pb) $\uparrow$ CLIP (pr) $\uparrow$ CLIP (pcam) $\uparrow$ CLIP (ph) $\uparrow$ FC (Within-Shot) $\uparrow$ FC (Cross-Shot) $\uparrow$ SC (Within-Shot) $\uparrow$ SC (Cross-Shot) $\uparrow$
EasyAnimate [69] 0.4086 0.2429 0.1633 0.1130 0.0722 0.4705 0.0268 0.7969 0.2037
CogVideo [71] 0.4113 0.2432 0.1632 0.1122 0.0701 0.6099 0.0222 0.7424 0.2069
VideoCrafter1 [8] 0.4365 0.2417 0.1535 0.1032 0.0651 0.3706 0.0350 0.7623 0.1867
VideoCrafter2 [9] 0.4000 0.2511 0.1654 0.1140 0.0694 0.5569 0.0686 0.7981 0.1798
VGoT 0.4086 0.2429 0.1633 0.1130 0.0722 0.8138 0.2688 0.9717 0.4276
*FC is denoted as Face Consistency, and SC is denoted as Style Consistency
Table 2: Human Evaluation. We compare VGoT with baseline models in terms of Within-Shot Consistency, Cross-Shot Consistency, and Visual Quality.
Within-Shot Consistency Cross-Shot Consistency Visual Quality
Bad $\downarrow$ Normal $\sim$ Good $\uparrow$ Bad $\downarrow$ Normal $\sim$ Good $\uparrow$ Bad $\downarrow$ Normal $\sim$ Good $\uparrow$
EasyAnimate [69] 0.3333 0.3232 0.3434 0.3535 0.3535 0.3131 0.4646 0.2727 0.2828
CogVideo [71] 0.1341 0.4146 0.4512 0.2927 0.5976 0.2317 0.1463 0.4512 0.5244
VideoCrafter1 [8] 0.5446 0.2574 0.1980 0.6436 0.1881 0.1683 0.6535 0.1782 0.1683
VideoCrafter2 [9] 0.1262 0.4854 0.3883 0.3495 0.3786 0.2718 0.1748 0.4951 0.3981
VGoT 0.0889 0.2556 0.6556 0.0889 0.2444 0.6667 0.0889 0.2111 0.7000
4.4 Module 4: Smooth Mechanism
The smoothing mechanism ensures a seamless transition from the current shot $i$ to the next shot $i+1$, ultimately achieving overall coherence throughout the final video. Inspired by FIFO [33] that uses latent partitioning to divide latents $Z$ into several parts to support different noise levels across frames, we implement a simple but efficient strategy to secure cross-shot video generation. Since adjacent shots could be dissimilar, we consider it necessary to reset the noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ for $\{z^{f}\}_{f=k\times i}^{(k\times(i+1)}$ generated with Eq 6 at shot $i$, and the length of the reset boundary should be equal to the frame number $k$ for each shot. Each reset boundary contains the conditional embedding $e^{T}_{i}$ and $e^{I}_{i}$, and a series of noise $\epsilon$ set with time step scheduler $\beta_{i}$ as defined in Sec 3. Until the final unit is finished, we can decode the whole $\{z^{f}\}_{f=0}^{F-1}$ into a final video $V$ with $N$ shots, and a visual demonstration of our generated results is shown in Fig 4.
Table 3: Ablation Studies. We evaluate the impact of removing key modules from our proposed framework. Metrics include CLIP Score, PSNR, IS, FC score, and SC score
CLIP average $\uparrow$ PSNR $\uparrow$ IS $\uparrow$ FC (Within-Shot) $\uparrow$ FC (Cross-Shot) $\uparrow$ SC (Within-Shot) $\uparrow$ SC (Cross-Shot) $\uparrow$
w/o EP w/o IP 0.1146 24.3265 7.4624 0.7364 0.1129 0.9406 0.3650
w EP w/o IP 0.1146 24.3265 7.5783 0.7305 0.1174 0.9471 0.3663
w/o EP w IP 0.1223 23.9228 7.4521 0.8745 0.3291 0.9486 0.4186
Full Model 0.1111 25.7857 7.5194 0.8303 0.2738 0.9487 0.3859
*FC is denoted as Face Consistency, and SC is denoted as Style Consistency
5 Experiments
5.1 Experiment Settings
Datasets
Since current datasets can’t satisfy our requirements for cross-shot story, as defined in Sec 3.2, we use VGoT to create 10 stories, each containing 30 shots, with a user input $S$ and 30 short shot scripts $S^{\prime}$, and a detailed 30 shot scripts $P$ from five domains as Eq 4, making a total of 300 video shots for comparison.
Metrics
In this work, we mainly utilize three metrics for quantitative evaluation: CLIP score [47], PSNR [17], IS [5], Face Consistency (FC) score and Style Consistency (SC) score. FC score is calculated by similarities between face features captured by InsightFace [48, 24, 21, 3, 2, 13, 14, 23, 11, 12] in generated videos, and we obtain SC score by style feature similarities from VGG19 [52]. Specifically, we calculate FC score and SC score from both within one shot and cross shots to evaluate the face identity and visual style consistency.
Implementation Details
Our experiments were conducted on A800 machines. For module implementation, we employed GPT-4o [1] as LLM, Kolor [59] as text-to-image diffusion model $\mathcal{M}_{I}$, DynamiCrafter [68] as conditional video model $\mathcal{M}_{V}$. For comparison, we also implement EasyAnimate [69], CogVideo [30], VideoCrafter1 [8],VideoCrafter2 [9].
5.2 Quantitative Evaluation
In this section, we use the 10 stories defined in Sec 5.1 as dataset, and provide the one-sentence $S$ for VGoT and the detailed scripts of 30 shots for the baselines as input. Then we use the 300 shots for each method to calculate CLIP score from five domains, together with FC score and SC score within and across shots. The quantitative comparison is shown in Table 1. Our method significantly outperforms the baselines in terms of both FC and SC, particularly when evaluated across shots, demonstrating that VGoT effectively maintains character identity and style consistency throughout the entire video, providing seamless transitions. It is also worth noting that the CLIP Scores for our method are not obviously comparable with the baselines. This phenomena is due to the diversity of the generated results, as evidenced by our ablation studies in Sec 5.5, which indicate that VGoT tends to produce richer, more varied content that may not strictly match the prompt but adds to the overall narrative depth and coherence.
5.3 Qualitative Evaluation
In this section, we use the one-sentence user input $S$ for VGoT, and detailed prompts $P$ generated by Script Module as input for baselines, to compare the visual effect of generated video results. Here we take the story “A set of one-sentence prompts, 30 shots, describe a story of a classic American woman Mary‘s life, from birth to death.” as an example, the comparison of results is demonstrated in Fig 3. Our method obviously achieves better visual consistency from style to identity over baselines.
5.4 Human Evaluation
In this section, we present a comprehensive user study comparing the generated results from VGoT and four baseline models. We use 10 stories, each consisting of 30 shots, as input and provide 10 users with a total of 50 multi-shot videos. Each user is shown 10 multi-shot videos selected randomly from the generated content and asked to evaluate them based on three criteria: within-shot consistency, cross-shot consistency, and visual quality, using a rating scale of good, normal, or bad.
The results, presented in Table 2, demonstrate that the videos produced by VGoT are preferred by users, particularly excelling in cross-shot consistency, highlighting the superiority of our method as acknowledged through human evaluation.
5.5 Ablation Studies
In this section, we present ablation studies to evaluate the impact of two key components: (1) Enhanced Prompt (EP) Generation, and (2) Identity-Preserving (IP) Control for keyframe consistency. Both quantitative and qualitative comparisons of VGoT are conducted by systematically removing either EP or IP modules, or both, using the user input: A set of one-sentence prompts, 30 shots, describing the journey of Carlos, from a young boy on his first bike to becoming a celebrated world champion cyclist.
The qualitative results, shown in Figure 5, illustrate that the complete VGoT framework (top row) maintains both character identity and narrative depth, while systematically removing EP (second row), IP (third row), or both (bottom row) results in a noticeable decline in visual diversity, narrative coherence, and character consistency.
The quantitative results, as presented in Table 3, further support our visual findings, with the highest PSNR and IS values indicating superior quality for our complete model. Two key insights from Table 3 are worth noting:
1. The full model achieves the lowest CLIP Score across all ablations. This is because the EP module enriches textual descriptions and the IP module preserves consistent character portrayal, resulting in generated visuals that deviate from the original prompts but ultimately enhance the video quality.
2. The removal of EP leads to improvements in FC and SC scores. However, while these metrics reflect increased cross-shot consistency, they do not capture the overall logic and diversity of the storyline. Comparing the middle and right portions of the top and second rows in Figure 5 shows that the absence of EP results in diminished narrative diversity, with visually similar shots benefiting consistency metrics but compromising the story’s richness. These ablation studies clearly demonstrate the importance of each component in producing high-quality, logically sound, and consistent multi-shot videos.
Figure 5: Visual Demonstration of the ablation studies of VGoT
6 Conclusion
In this paper, we introduce VideoGen-of-Thought (VGoT), a modular framework that tackles the challenges of generating multi-shot videos with narrative coherence, visual consistency, and stylistic fidelity. VGoT breaks down video generation into four collaborative modules, ensuring high semantic alignment, keyframe consistency, smooth transitions, and scene-specific guidance. Our evaluations demonstrate that VGoT outperforms existing baselines in maintaining flow consistency, face coherence, and stylistic alignment, offering a promising approach for scalable and coherent multi-shot video generation.
Supplementary Material
Appendix A Implementation Details of Script Module
The Script Module plays a fundamental role in converting high-level user input into a detailed series of prompts for each shot within the multi-shot video generation process. The specific process is to convert a single sentence user input $S$, into a more detailed and structured description $S^{\prime}$, which is then decomposed into a set of prompts $P=\{p_{1},p_{2},\dots,p_{N}\}$, corresponding to each of the $N$ shots required for the complete video. This process uniformly adopts a large language model (LLM) and uses prompt engineering to ensure the reasonableness of the generated video. This process enables the generation of a logical, stepwise narrative structure that serves as the backbone for subsequent keyframe and shot-level generation.
For example, consider the user input: e.g., a set of one-sentence prompts, 30 shots, describe a story about three avatars (two friend Mike and Jane, one bad guy Tom) and their fantasy adventure in ancient pyramid.. The LLM would first transform this input into a detailed version $S^{\prime}$ consisting of 30 one-sentence scripts like $s_{1}$: “Mike’s Discovery: Mike examines an ancient map with intense focus, revealing the path to the pyramid.” Subsequently, each $s_{i}$ in $S^{\prime}$ is decomposed into prompts $P$, such as $p_{1}$: Character: Mike, holding an ancient map with Jane by his side. Background: A dense jungle filled with mist and towering trees. Relation: Mike studies the map closely, pointing to a pyramid. Camera Pose: Medium shot focusing on Mike and Jane. HDR Description: Soft light filters through the trees, creating dynamic shadows on the map and characters.”. This structured approach ensures narrative coherence across multiple shots, laying the foundation for the generation process.
Appendix B Additional Results
To further illustrate the effectiveness of our proposed VideoGen-of-Thought (VGoT) framework, we present additional qualitative results showcasing the generated videos. Besides the pyramid story with three avatars provided in Sec A and materials, VGoT achieves a better reasonable storyline and visual consistency in stories about one protagonist’s growth at different times. We prepared ten showcases in the materials, and a large visual representation of VGoT’s output is provided in Figure 6, spanning two full pages to ensure sufficient detail and clarity for each generated shot. This figure demonstrates the stepwise progression of narrative shots, showcasing the seamless visual consistency and logical coherence in a wide variety of generated multi-shot videos.
Moreover, we also provide additional comparison results to illustrate VGoT’s advantages over existing state-of-the-art methods. These comparisons include visual comparisons with four baselines: EasyAnimate, CogVideo, VideoCrafter1, and VideoCrafter2. Each comparative example is analyzed in terms of visual consistency, narrative coherence, and overall quality. As shown in Figure 7, VGoT consistently outperforms the baselines in terms of character continuity, background stability, and logical flow across shots. These results highlight VGoT’s ability to maintain coherent storytelling while also achieving high-quality visuals. We also prepared the original experiment data record for quantitative evaluation and ablation studies in our provided materials.
Appendix C User Study
To evaluate the user-perceived quality of videos generated by our VGoT framework, we conducted an extensive user study involving 10 participants. The participants were given 50 accelerated multi-shot videos, each generated either by VGoT or one of four baseline methods. The 10 input stories, consisting of 30 shots each, were randomly assigned to ensure diverse feedback and minimize bias. Each user was presented with 10 videos from different sources and asked to evaluate them on a scale of good, normal, or bad, based on three specific criteria: within-shot consistency, cross-shot consistency, and visual quality.
The results of the user study are summarized in Figure 8. The data indicates that users significantly preferred the videos generated by VGoT, especially regarding cross-shot consistency. Users found that VGoT’s videos maintained logical transitions between shots and preserved character appearances across different scenes, reflecting the robustness of our approach. Compared to the baselines, VGoT’s results were rated highly for narrative coherence and overall quality, demonstrating the effectiveness of our collaborative multi-shot framework in meeting user preferences.
Appendix D Limitations and Future Work
While our VGoT framework offers significant advancements in multi-shot video generation, there are still limitations that need to be addressed, along with directions for future work.
1. Single Identity-Preserving Embedding per Shot: Currently, each shot only receives one identity-preserving (IP) embedding, which limits the flexibility to depict complex scenarios involving multiple characters or subjects. Future work will explore extending the framework to handle multi-subject driven shot-level video generation, thereby enhancing narrative richness and visual diversity.
2. Limited Evaluation Metrics: Our current evaluation focuses mainly on face and style consistency, which may not capture all aspects of multi-shot video quality comprehensively. There is a lack of detailed multi-shot evaluation metrics to assess narrative flow and cross-shot coherence. Future work will involve researching advanced cross-shot metrics and using these metrics to inform and guide model training, potentially improving overall video generation quality.
Figure 6: VGoT Visual complement of the multi-camera video generated.
Figure 7: Visual comparison of VGoT with baselines Supplement.
Figure 8: Designed user study interface. Each participant was required to rate 50 videos by answering three sub-questions for each video. Due to page limitations, only two videos are shown here.